OpenStudy (anonymous):

Help? I have to graph this, then find domain, range, x and y intercepts, and horizontal and vertical asymptotes.

OpenStudy (anonymous):

\[\frac{ 2 }{ x^2-2x-3}\]

OpenStudy (anonymous):

I believe the Joker knows the answer :DDD

OpenStudy (anonymous):

Rude. :p

OpenStudy (anonymous):

Whoa.. None of the equation things are working.. Maybe it's my computer..

OpenStudy (amoodarya):

x2−2x−3=(x-3)(x+1)

OpenStudy (anonymous):

\[\frac{ 2 }{ (x-3)(x+1) }\] the domain is \[\mathbb{R}-\left\{ -1,3 \right\}\]

OpenStudy (amoodarya):